Two small spheres spaced apart have equal charge. How many electrons must be present on each sphere if the magnitude of the force of repulsion between them is 4.57 x 10^-21 N?


1
Expert's answer
2022-04-07T18:25:46-0400

"F=\frac{kn^2e^2}{r^2},"

"n=\sqrt \frac Fk\frac re=400."